A gas canister can tolerate internal pressures up to 159,600 mmHg. If a 2.0 L canister holding 3.5 moles of a gas is heated to 1350⁰C, will the canister explode?
1
Expert's answer
2020-02-07T05:28:46-0500

Mendeleev Clapeyron's Law

nRT=PV

n - the number of moles, R - gas conct. 8,31 J/moles*K=0,0821l*atm/mole*K, T - temperature (*K=*C+273,15), P - pressure Pa, V- volume of ideal gas, L,

Given: n=3,5 moles, T=1350+273,15=1623,15, P max=159600 Torr=210 atm= 21278250Pa, V=2.0 litr

Reality: P=nRT/V=~233,206 atm

Preal>Pmax


canister explode